OpenStudy (anonymous):

can you help me with this for my exercises: I need to to show that: g(x)= (e(-3x+2sinx)/6) ---------------- (3+2cosx) has derivative g'(x)= - (4(sin^2)x -12sin x +5) (e(-3x+2sinx)/6) ----------------------------------- 6(3+2cosx)^2

OpenStudy (anonymous):

\[d(u/v)=(v*u'-v'*u)/v^2\] try this

OpenStudy (chaise):

The above is right, but also try re-write your equation like this: then applying what nageswar has used (the quotient rule) \[\frac{e(2\sin(x)-3x)}{6(2\cos(x)+3)}\]. I think you will find that dy/dx in this equation is not the equation given (the derivative of sin(x) is cos(x) which has not changed in the equation at all. Unless you have written the question wrong or I am misinterpretting it.